OpenStudy (anonymous):

9y^2-6y-9-x=0 complete the square and reduce to one standard form y-b=A(x-a)^2 or x-a=A(y-b)^2

OpenStudy (anonymous):

\[9y^2-6y-9-x=0\] \[9y^2-6y=x+9\] \[9(y^2-\frac{2}{3}y)=x+9\] \[9(y-\frac{1}{3})^2=x+9+9\times \frac{1}{9}\] \[9(y-\frac{1}{3})^2=x+10\]